OpenStudy (anonymous):

solve for u: du/dt = 2(5-u).

OpenStudy (watchmath):

\(\int du/(5-u)=\int 2 dt\) \(-\ln|5-u|=2t+C\) \(\ln|5-u|=-2t+C\) \(|5-u|=Ae^{-2t}\) (where \(A\) is just another constant) \(5-u=\pm Ae^{-2t}\) \(u=5\pm A e^{-2t}\)